A truck driver is driving at a velocity of 60 km/h. He sees a danger ahead and steps on his brakes. He stop the truck in exactly 5 seconds. How far does he go before coming to a stop?

1
Expert's answer
2021-12-16T11:32:16-0500

"v_0 = 60\frac{km}{h}=\frac{60*1000}{3600}\approx16.7\frac{m}{s}"

"v_1 = 0"

"t = 5s"

"a = \frac{v_1-v_0}{t}= -3.34\frac{m}{s^2}"

"s = v_0t+\frac{at^2}{2}=16.7*5- \frac{3.34*25}{2}= 41.75m"


"\text{Answer: } 41.75m"